### v1

# Competing effect of disorder on phase separation in active systems: from facilitated to hindered growth and rough interfaces

Abstract
We investigate the impact of random pinned disorder on a collection of self-propelled particles. To achieve this, we construct a continuum model by formulating the coupled hydrodynamic equations for slow variables: local density and momentum density of particles. The disorder in the system acts as pinning sites, effectively immobilizing the particles that come into contact with them. Our numerical results reveal that weak disorder leads to phase separation in the system at density and activity lower than the typical values for motility induced phase separation. We construct a phase diagram using numerical simulations as well as linearized approximation in the plane of activity and packing fraction of particles at weak disorder densities. As disorder densities rise in the system, kinetic processes slow down, while at high disorder densities, the system becomes heterogeneous and eventually undergoes kinetic arrest. The structure factor tail deviates from Porod’s law, indicating increased roughness at domain interfaces under strong disorder. Furthermore, we analyze the fractal dimension of the interface as a function of disorder density, highlighting the increasing irregularity of phase-separated domains.
I Introduction
Nonequilibrium living and lab designed active systems for example, cells, tissues, living organisms and autonomous robots Ramaswamy (2010, 2017); Marchetti et al. (2013); Vrugt and Wittkowski (2024); Volpe et al. (2022) composed of many self-propelled agents unveil intriguing collective behavior across a wide range of length and time scales. The study of active matter systems became an interesting area of research for many years due to the emergent behaviors such as pattern formation Denk and Frey (2020); Hallatschek et al. (2023); Jena and Mishra (2024), nonequilibrium disorder to order transitions Kumar et al. (2024), anomalous fluctuations Jena and Mishra (2023); Ramaswamy et al. (2003), and interesting behaviour in different medium and confinement etc. within these systems, that are not present in corresponding analogous equilibrium systems. Another captivating characteristic of such systems is Motility Induced Phase Separation (MIPS) that resembles the passive liquid-gas phase separation but occurs in absence of any attractive interactions Fily and Marchetti (2012); Cates and Tailleur (2015) at much lower packing densities.
Majority of studies of active matter in theory and experiments are focused on the systems in homogeneous or clean environment Jülicher et al. (2018); Sanchez et al. (2012); Zöttl and Stark (2016); Alston et al. (2022); Aranson (2022); Gachelin et al. (2014); Pattanayak and Mishra (2018); Sampat and Mishra (2021); Takatori and Brady (2016); Fily and Marchetti (2012); Wittkowski et al. (2014); Stenhammar et al. (2013). But, in natural active matter systems, inhomogeneity or disorder is present intrinsically. Various types of disorder or inhomogeneities are observed in natural systems that arise from multiple factors. This can be present in the form of spatial (geometric variation), temporal (moving obstacles), chemical composition and, biological inhomogeneity (bio-diversity) etc. Exploring and understanding such systems with disorder is essential to realize the complexity of natural systems and the systems’ response to internal as well as external stimuli.
There has been extensive studies on active systems in clean environment, while the exploration of systems with disorder is relatively less. However, there exist a few prior studies on systems possessing disorder Kumar and Mishra (2020); Das et al. (2018); Chen et al. (2022); Vahabli and Vicsek (2023); Pinçe et al. (2016); Ro et al. (2021); Goswami and Chakrabarti (2022); Mishra and Mishra (2023). These studies explore the effects of both quenched and annealed disorder in active systems, highlighting their impact on the dynamics of active particles by examining transport properties and diffusivity. In certain cases, disorder enhances collective behavior in active polar particles, while in active scalar particles, it has been reported to induce long-range order Ro et al. (2021). The findings Kumar et al. (2021); Sándor et al. (2017); Reichhardt and Reichhardt (2018) reveal that the disorder can lead to the emergence of distinct phases. In most studies, disorder acts as either obstacles placed physically in the space or a random field disorder Das et al. (2018); Chepizhko and Peruani (2013); Ro et al. (2021); Morin et al. (2017). Apart from some studies, very little is still known about the phase behavior of the active systems in the presence of disorder which acts like pinning sites and immobilize the particles. This kind of fixed obstacles can be noticed in many natural system i.e, motion of bacteria in porous media Drescher et al. (2011), migrating cells encountering collagen fibers Wolf and Friedl (2011), collective cell migration through tissue environment Petrie et al. (2009) etc.
Our primary focus in the present work is on the effect of pinned disorder on the phase separation and kinetics of collection of self-propelled particles, setting it apart from previous works. The study will be beneficial for observing the dynamical behavior of the organism in the presence of disorder in natural system.
In the present work, we introduce disorder, as pinning sites in a system consists of a collection of active particles, by constructing a continuum model by formulating the coupled hydrodynamic equations for slow variables: local density and momentum density of particles. The pinning sites are modeled in such a manner that they effectively make the self-propelled speed of the active particles zero. We explore the phase diagram to gain insight into the phenomenon of phase separation in the presence of disorder. We observed phase separation at density and activity lower than the critical values for a clean system Fily and Marchetti (2012); Redner et al. (2013), a result further confirmed by linearized hydrodynamic calculations. Although the pinned disorder eases the phase separation, the kinetics of phase separation slows down at lower disorder and becomes arrested at high disorder densities. The structure factor tail exhibits a deviation from Porod’s law, suggesting enhanced roughness at domain interfaces in the presence of strong disorder. Additionally, our analysis of the fractal dimension of the interface as a function of disorder density reveals irregularity in phase-separated domains.
The structure of this paper has been organized as follows. SecII provides the detail explanation of our model along with the numerical methodology for solving the equations. We demonstrate the outcomes of our research in secIII. In secIV, we discuss our interesting notable findings and present a summary highlighting the relevance of our study.
II Model and Numerical Methodology
We develop a continuum model to study a collection of self-propelled particles on a two dimensional substrate with the presence of random pinning sites. We formulate a minimal hydrodynamic coupled equations for conserved local density field $\rho(\bm{r},t)$ and local orientation order parameter $\bm{p}(\bm{r},t)$ and further we define the momentum density $\bm{P}(\bm{r},t)=\rho(\bm{r},t)\bm{p}(\bm{r},t)$. The hydrodynamic equation for the density $\rho(\bm{r},t)$ is:
$\partial_{t}\rho=-\bm{\nabla}.(v(\rho)\bm{P}-D_{\rho}\bm{\nabla}\rho+\bm{f}_{\rho})$ (1)
and for the local momentum density $\bm{P}(\bm{r},t)$ is:
$\partial_{t}\bm{P}=-\nu_{r}\bm{P}-\frac{1}{2}\bm{\nabla}(v(\rho)\rho)+k\bm{\nabla}^{2}\bm{P}+\bm{f_{P}}$ (2)
here, $v(\rho)=v_{0}(1-\lambda\rho)$ denoted as the effective velocity of each particle, is influenced by the local density of the particles and the presence of pinned sites in the system. Its dependence of local density is adopted from previous observations that the local clustering suppresses the motile nature of particles Fily and Marchetti (2012). Further, we introduce the pinned obstacles or disorder in such a manner there are few randomly selected regions in the substrate such that $v(\rho)=0$ at those points. The number density of such points on the substrate is the density of pinning sites in the system is defined as $\rho_{d}$.
Eq.1 is a continuity equation where, the first term on the right side of the equation accounts for the active current generated by self-propulsion and the second term represents the diffusion current with diffusion constant $D_{\rho}$. Eq. 2 is similar to the equation introduced by Fily et.al Fily and Marchetti (2012) in the context of self-propelled particles neglecting the nonlinearites in $\bm{P}$. The first and second terms on the right hand side of Eq. 2 represents the polarization decays at rate $\nu_{r}$ and is convected by pressure-like gradients $\sim\bm{\nabla}(v(\rho)\rho)$. The third term represents the diffusion in orientation. And $\bm{f}_{\rho}$ and $\bm{f}_{\bm{P}}$ are the Gaussian white noise in density and polarization equation respectively having strengths $\Delta_{\rho,\bm{P}}$;
$\langle f_{\alpha,i}({\bf r},t)f_{\alpha,j}({\bf r^{{}^{\prime}}},t{{}^{\prime}})\rangle=\Delta_{\alpha}\delta_{ij}\delta({\bf r}-{\bf r}^{{}^{\prime}})\delta(t-t^{{}^{\prime}})$
where, $\alpha\equiv(\rho,\bm{P})$ and $(i,j)$ can take two values $x$ and $y$ representing the two cartesian coordinates. The intrinsic time scale $\tau$ and intrinsic length scale $l_{0}$ are defined as $\nu_{r}^{-1}$ and $\sqrt{D_{\rho}/\nu_{r}}$ respectively. The Eq. 1, 2 are rescaled by the $\tau$ and $l_{0}$. Numerical integration of Eq. 1 and 2 is performed with homogeneous initial density with mean $\rho_{0}$ and random ${\bm{P}}$. We vary the mean density from $\rho_{0}$ = 0.2 to 1.0 and self-propulsion speed from $v_{0}=0$ to $7$ in a box of size $K\times K$ with periodic boundary condition in the both directions. To investigate the impact of disorder in the system, we vary the disorder density $\rho_{d}$ from 00 to $0.5$. The integration is performed using Euler’s scheme Bally and Talay (1996) with $\Delta x=1.0l_{0}$ and $\Delta t=0.1\tau$. In the system under consideration, the parameters are fixed as follows: $\nu_{r}=k=D_{\rho}=1.0$ and $\lambda=0.9$. For these choice of parameters the initial homogeneous state becomes unstable and system phase separates for $\rho_{0}>1/(2\lambda)$ as reported in Fily and Marchetti (2012). We performed numerical simulations for simulation time, $t=2\times 10^{5}$ and for system sizes $K=256-1024$ and averaging is performed over $50$ independent realizations.
Figure 1: The panels (a-d) showcase the local fluctuation $\delta\rho$ at across four panels , each representing system with different disorder densities $\rho_{d}=0.0,0.05,0.1,0.5$ respectively at $t=200$. Within each panel, multiple figures are plotted from left to right for self-propulsion speed $v_{0}=0.2,1,3,5,7$ keeping the mean density $\rho_{0}=0.6$ fixed. The color on the heatmap represents the value of $\delta\rho$. The results are obtained for systems having $K=256$.
Figure 2: The panels (a-d) depict the local fluctuations in density, $\delta\rho$, at time $t=200$ for systems with varying disorder densities, $\rho_{d}=0,0.05,0.1,0.5$ in sequence. Each panel consists of multiple snapshots, from left to right, representing systems with mean densities $\rho_{0}=0.2,0.4,0.6,0.8,0.9$ respectively by fixing the self-propulsion speed $v_{0}=4.0$. The color in the heatmap indicates the magnitude of the $\delta\rho$. The results are generated for systems having $K=256$.
Figure 3: The plot(a) presents semi-log $x-$ plot of $\Delta\rho$ vs. $\rho_{d}$ with error bars. The inset shows a linear plot of $\Delta\rho$ vs. $\rho_{d}$ depicting the non-monotonicity with respect to disorder. The plot(b) illustrates the phase diagram in the $\rho_{0}v_{0}^{2}-v_{0}^{2}$ plane for different disorder densities: $\rho_{d}=0.0,0.001$ respectively. The blue (circles) and red (squares) solid lines are the boundary drawn from the numerical simulation for $\rho_{d}=0.0,0.001$ which is also mentioned in the legend. Additionally, the green solid line and magenta dotted line represent the analytical boundary obtained from linearized calculation. The shaded light-orange region shows the extra regime of phase separation in the presence of disorder. The plot (c) showcases $\Delta\rho$ vs. $v_{0}$ depicting the transition from non-phase separation to phase separation for $\rho_{d}=0.0,0.001,0.01$.
Figure 4: The panels (a-d) showcase the snapshots of time series of the local density field $\rho$ at t = 5, 50, 150 and 200 across the columns in sequence. Each panel from top to bottom, depicts the snapshots for system with $\rho_{d}=0,0.05,0.3,0.5$, respectively. The color in the heatmap represent the magnitude of local density at each lattice point. The results are obtained from simulating a $256\times 256$ system.
III Result
Snapshots for local density fluctuation:
To investigate the effect of disorder on the system, we begin by plotting snapshots of normalized local density fluctuations (${\delta\rho}$) defined as ${\delta\rho}({\bm{r}},t)=\frac{\rho({\bm{r}},t)-\rho_{0}}{\rho_{0}}$. We examine the result for a range of ($v_{0}$, $\rho_{0}$) values in systems with varying density of disorder $\rho_{d}$. Figs. 1(a-d) illustrate snapshots of local densities fluctuations across four panels, showcasing the effects of varying $\rho_{d}$ and fixing $\rho_{0}=0.6$. Panel (a) represents a system without disorder ($\rho_{d}=0$) for different values of $v_{0}=0.2,1,3,5,7$. Panels (b), (c), and (d) display the corresponding results for systems with disorder densities of $\rho_{d}$ = 0.05, 0.1, and 0.5, respectively. Fig. 1(a) shows that in a clean system, phase separation begins at $v_{0}>3$ with the formation of small domains, indicating the onset of Motility-Induced Phase Separation $(MIPS)$, consistent with previous studies on active Brownian particles Fily and Marchetti (2012). For the systems with finite density of disorder, ${\delta\rho}$ starts to develop finite values at $v_{0}=3$ as can be seen in Fig. 1(b) for $\rho_{d}=0.05$. Thus, the density inhomogeneity develops at a speed below that for the clean system. As the numbers of pinning sites increases, as shown in Fig 1(c) for $\rho_{d}=0.1$, the density inhomogeneity starts to appear at lower $v_{0}$. But, after a certain $\rho_{d}$ the system does not phase separate and instead shows the heterogeneous density as shown in Fig. 1(d) for $\rho_{d}=0.5$. Similarly in Fig. 2(a-d), we show the fluctuations in local density for fixed $v_{0}=6$ , varying mean particle densities $\rho_{0}$ , for different disorder densities.
The density of disorder is the same as for the Fig. 1(a-d). This allows us to examine how both $\rho_{0}$ and disorder influence the system’s behavior. We observed that the system undergoes phase separation at ${\rho_{0}}\geq 0.6$ for the clean system ($\rho_{d}=0$) in Fig. 2(a). As suggested by previous studies for clean systems Fily and Marchetti (2012), $MIPS$ occurs once a critical density is reached. However, as the disorder in the system increases, phase separation occurs even at lower values of $\rho_{0}$ as shown in Fig.2(b-c). Therefore, the phase diagram in the ${\rho_{0}}-\rho_{d}$ plane demonstrates that disorder promotes density inhomogeneity in the system for weak disorder. Similar to Fig.1(d), higher disorder makes system heterogeneous again as shown in Fig.2(d).
Based on the observation from the snapshots shown in Figs. 1 and 2, it can be concluded that the system with finite $\rho_{d}$ shows the phase separating domains at self propulsion speed $v_{0}$ and mean density $\rho_{0}$ smaller than that for the clean system $\rho_{d}=0$. Hence the presence of small number of pinning sites drives the phase separation at lower activity and mean density. Further, the system with high density of disorder shows the formation of heterogeneous structures suppressing phase separation.
To analyze it thoroughly, in Fig. 3(a) we show the plot of density phase separation order parameter $\Delta\rho$ (PSOP) vs. $\rho_{d}$, where $\Delta\rho=\frac{1}{K^{2}}<{\sum}_{\bm{r}}|\delta\rho(\bm{r},t)|>$, where <..><..>< . . > mean average over time in the steady state and over different realizations. $N$ is the number of lattice points in the system, $\Delta\rho$ reflects the amount of phase separation in the system. We find $\Delta\rho$ shows the non-monotonic feature as we increase the disorder density. We fix the self-propelled speed $v_{0}$ and $\rho_{0}$ to values 4.0 and 0.6 respectively, such that the homogeneous state is stable so that $\Delta\rho=0$ for the clean system. As we introduce disorder, $\Delta\rho$ shows a jump to finite value and again decreases for the large disorder densities as shown in Fig. 3(a)(inset). A very small disorder is enough to make the homogeneous state unstable as shown in Fig. 3(a)(main panel) (on semi-$\log x$- scale).
We now investigate the mechanisms responsible for disorder-induced phase separation at low disorder densities and the transition to a heterogeneous phase at high disorder densities. In the low disordered environment, the accumulation around the obstacles could be a result of particles being immobilized upon reaching close proximity, which leads to effectively zero velocity causing the particles to gather in that location. Additionally, when new particles approach, the existing ones act as obstacles and contributing to the reduction of their velocity. As a result of this effect, clustering around the obstacles and domains formation is observed. Since, the phase separation is driven by the disorder in the system, we refer to it as Disorder-Induced Phase Separation ($DIPS$). However, for high $\rho_{d}$, the particles get immobilized at most of the sites. There is no scope for movement of the particles and they get stuck immediately around their nearest pinning sites. As the particles immobilize immediately, they can not accumulate in sufficient amount to form large clusters. Although the diffusion in density term is non-zero, but it will simply diffuse the particles from one place to another and pinned sites will act like cold regions with small accumulation of density around that sites. As a result, heterogeneous structures are formed in the system.
Phase Diagram: To better illustrate the shift in the phase boundary, we present a phase diagram in Fig.3(b). To gain further insight from the hydrodynamic equations, we perform a linearized analysis around the homogeneous state, demonstrating that in the activity($v_{0}$) and mean density ($\rho_{o}$) plane, the phase boundary shifts towards lower values of both.
We write the linearised equations for the small fluctuations around the homogeneous phase; i.e., ${\bm{P}}=\delta\bm{P}$ and $\rho=\rho_{0}+\delta\rho$. The details of the calculation is shown in Appendix. Using the condition of instability, we find a relation between activity $v_{0}$ and the mean density $\rho_{0}$ where the homogeneous state becomes unstable as given in Eq. A10 in the appendix. In the Fig. 3(b), we show the linear relation between $\rho_{0}v_{0}^{2}$ vs. $v_{0}^{2}$ for the two cases: system without disorder $\rho_{d}=0$ and system with disorder $\rho_{d}=0.01$. For system with $\rho_{d}=0$, in the region above the solid line, the homogeneous phase is unstable, whereas in the presence of finite disorder $\rho_{d}=0.01$, region above the dashed line is also unstable. Hence, the additional area shown by shaded color is the parameter space, where the homogeneous state becomes unstable due to the presence of disorder. This observation is consistent with our numerical simulations, where the points in Fig. 3(b) are obtained from simulation data. Consequently, in the presence of weak disorder, the phase boundary shifts to lower values, indicating that disorder facilitates phase separation at reduced mean density and particle activity.
We further investigate the transition from homogeneous to phase separated state for different disorder densities and activities. In Fig. 3(c) we show the plot of phase separation order parameter $\Delta\rho$ vs. $v_{0}$ for different: $\rho_{d}=0.0$ $0.2$ by fixing $\rho_{0}=0.6$. The $\Delta\rho$ shows a jump at $v_{0}\sim 5.0$ for the clean system, in contrast in the presence of disorder, the change in $\Delta\rho$ happens at lower $v_{0}\sim 3.0$ and slowly on increasing disorder the $\Delta\rho$ continuously changes from zero to finite values. Accordingly, the disorder makes the transition continuous as found in previous studies Villa Martín et al. (2014).
So far, we discussed the role of disorder on the steady state of the system. Next, it will be interesting to understand how the presence of disorder affects the kinetic of phase separation and morphology of domain walls. To do so, we focus on the parameter space where the homogeneous state is inherently unstable and undergoes phase separation.
Figure 5: The main plot (a-e) present the scaled $g(x)$ vs. $x$ for disorder density $\rho_{d}=0-0.5$ respectively. The insets of each plot shows time progression ($t=1000-20000$) of correlations by plotting the $C(r,t)$ vs. $r$ for corresponding disorder density. The correlation functions are obtained from system of $K=512$ averaging over 50 ensembles. The snapshots (f-j) showcase the local density fluctuation $\delta\rho$ at $t=1000$ for $\rho_{d}=0.0,0.05,0.1,0.3,0.5$. The heatmap shows the magnitude of $\delta\rho$ at each lattice point. We have taken a $256\times 256$ system to obtain the snapshots.
Figure 6: The plot (a) showcases the log-log plot of characteristic length $L(t)$ vs. $t$ for different disorder density $\rho_{d}$ ranging from $0-0.5$ in the system showing in the legends. The dashed orange line depicts the line of slope $0.33$. The plot (b) displays semi-log $x-$ plot of the corresponding $1/z_{eff}$ with $\rho_{d}$. The error bars represent the standard deviations of the $1/z_{eff}$. The plot (c) presents the static scaled correlation $g(x)$ for $\rho_{d}=0-0.5$.
Kinetics of phase separation: Fig. 4 (a-d) depict the snapshots of the local density fluctuations $\delta\rho(\bm{r},t)$ in the four panels for $\rho_{d}=0-0.5$ in sequence at different times $t=5-200$ as shown in each row of the panels. From the snapshots, it can be seen that the density inhomogeneity is observed for $\rho_{d}=0,0.05,0.3$ in Fig. 4(a-c). However, in case of higher disorder i.e, $\rho_{d}=0.5$ , the density homogeneity is suppressed showing almost no phase separation in Fig. 4(d). It would be insightful to examine the growth law for domains in a disordered system compared to a clean system. A standard tool to obtain information about sizes and textures of domains and interfaces is the two-point spatial correlation function of density fluctuations defines as $C(r,t)$ = $\langle\delta\rho({\bf r}^{\prime}+{\bf r},t)\delta\rho({\bf r}^{\prime},t)\rangle$ and the corresponding Fourier transform i.e, structure factor $S(k,t)=\langle\delta\rho({\bf k}^{\prime}+{\bf k},t)\delta\rho({\bf k}^{\prime},t)\rangle$. The $\langle$...$\rangle$ … . ⟩ denotes an average over reference positions ${\bf r}^{\prime}$, and $50$ independent realizations. In Fig. 5(a-e), we present the correlation function for system having disorder densities ranging from $\rho_{d}=0-0.5$ respectively. The insets in each plot depict the progression in time of the correlation function. The decay of correlations is slower with time, suggesting an increase in the size of clusters. The time evolution of the correlations are similar in the system with $\rho_{d}=0,0.05,0.1$ in Fig. 5(a-c) (inset), while a slower increase in correlations observed in system with $\rho_{d}=0.3$ in Fig. 5(d) (inset). Further, in case of high disorder density $\rho_{d}=0.5$, the temporal development of correlation is arrested showing sharp decay in correlations as shown in Fig. 5(e) (inset).
The main plots of Fig. 5(a-e) display the scaled correlation functions obtained as $g(x)$ , where $x=r/L(t)$. Here, $L(t)$ is the characteristic length determined as the distance over which the correlations crosses 0.1. The correlation functions show a dynamical scaling showing a good scaling collapse across all disorder cases. That implies the existence of a single characteristic length scale $L(t)$ and the evolution morphology can be characterized by a distinct single length scale for each disorder level. Although, the weak disorder promotes phase separation, the decrement in the two-point correlation functions in the disordered system, make it distinct from the $MIPS$. In Fig. 5(f-j), we show the snapshots for local density fluctuation $\delta\rho(\bm{r})$ in the system having the corresponding values of $\rho_{d}=0-0.5$ respectively at time $t=1000$. In Fig. 5(f-h), we observe domain formation showing stronger density contrast up to $\rho_{d}=0.1$, while in the case of $\rho_{d}=0.3$, phase separation is observed with smaller density inhomogeneity as shown in Fig. 5(i) and there is formation of heterogeneous structures and no clear phase separation for $\rho_{d}=0.5$ as shown in Fig. 5(j). The similar results are obtained in previous study of Ro et al. (2021) for disordered system. These snapshots justify the above correlation plots nicely.
Next, to illustrate the impact of disorder in the growth law, in Fig. 6(a) we present a log-log plot of $L(t)$ vs. $t$ for systems with varying levels of disorder $\rho_{d}$ from $0.0$ to $0.5$. The plot clearly shows that, asymptotically, the slope of $L(t)$ for $\rho_{d}=0.0,0.05,0.1$ is approximately $0.33$, consistent with the usual Lifshitz-Slyozov (LS) growth law Bray (2002) suggesting that phase separation in clean and low-disorder systems follow the standard LS law. However, as the disorder increases further i.e, for $\rho_{d}=0.3,0.4$, the slope gradually decreases and signifying the deviation from $LS$ growth law for systems with higher disorder. Further, at very high disorder levels, such as $\rho_{d}=0.5$, the plot shows no significant growth of $L(t)$, indicating a complete suppression of the usual phase separation kinetics.
To interpret this more quantitatively, we estimate the effective growth exponent $1/z_{eff}$ as a function of $t$ defines as $\frac{1}{z_{eff}}=\bigg{<}\frac{dlnL(t)}{dlnt}\bigg{>}$, where <..><..>< . . > means average over late times. In Fig. 6(b), we showcase a semi-log $x$ plot depicting the variation of $\frac{1}{z_{eff}}$ with $\rho_{d}$. The plot represents that the dynamic growth exponent remains constant at 0.33 up to $\rho_{d}=0.1$, but gradually decreases towards zero as the disorder increases. This suggests a suppression of domain growth with increasing disorder in the system.
In Fig. 6(c) we show the scaled plot of $g(x)$ vs. scaled distance $x$ for different disorder densities, same as in Fig. 5. Very clearly for $\rho_{d}\leq 0.1$, all the correlations show the nice collapse, indicating the static scaling for lower disorder densities, whereas the static scaling is not found for $\rho_{d}>0.1$.
Figure 7: The plot (a) exhibits the plot of scaled structure factor $S(k)L^{-2}$ vs. $kL$ for $\rho_{d}=0-0.5$ as shown in the legend. The maroon and orange dotted lines present the slope of line $-3$ and $-2.5$. The plot (b) displays $1-g(x)$ vs. $x$ for different disorder density and the symbols have the same meaning as shown in (a). The legends present the slope of lines. (c) showcases the log-log plot of $d_{f}$ vs. $\rho_{d}$. The error bars are of the size of the symbols used. The plot (d) depicts the density fluctuation $\Delta\rho_{n}$ vs. $\rho_{n}$ for different $\rho_{d}$. The symbols have the same meaning as in plot (a). The legends show the slope of lines as mentioned.
Domain morphology and density fluctuations:-
Till now we have focused on the growth kinetics but disorder also affects the domain morphology. To analyze the morphology in more details, including the interfacial properties, we calculate the scaled structure factor $S(k)L^{-d}$ and the cusp exponent $1-g(x)\sim x^{\theta}$. The exponent $\theta$ depends on the relevant morphology and reveals the roughness of the interface. In Fig. 7(a), we present the plot of scaled structure factor $S(k)L^{-2}$ vs. $kL$ for system having $\rho_{d}$ from $0.0$ to $0.5$. Fig. 7(a) are already scaled with respect to time for individual system having different $\rho_{d}$. The range of time is the same as that for the $C(r)$ shown in Fig. 5(a-e). We found all the curves for $\rho_{d}\leq 0.1$ shows the nice static scaling as well, however, static scaling is not observed for $\rho_{d}>0.1$ as shown in Fig. 7(a). It suggest that the morphologies of the domain changes on varying disorder densities. For disorder densities up to $0.1$, we observe the Porod’s tail $S(k)\sim k^{-3}$ implying the sharp interface between the domains. However, above $\rho_{d}>0.1$, the behavior deviates from Porod’s tail, approaching $S(k)\sim k^{-2.5}$ which suggests the emergence of indistinct boundary between domains. We also compared the structure factor calculated from linearised hydrodynamic and it shows the structure factor flattens with increasing disorder the same as found in the numerically in Fig. 7(a).
Further, to examine the roughness of the interface, in Fig. 7(b), we plot $1-g(x)$ vs. $x$ for different disorder densities ranging from $\rho_{d}=0.0-0.5$. We observe a crossover of $\theta$ from 1 to 0.5. For $\rho_{d}\leq 0.1$, the morphology of the domains are smooth, whereas we notice the rough domains for $\rho_{d}>0.1$. We examine the fractal nature of the interfaces by calculating their fractal dimension using the box-counting method. The detailed method is provided in the appendix. We showcase a log-log plot of the fractal dimension of the interface $d_{f}$ as a function of disorder density $\rho_{d}$ in Fig. 7(c). The plot reveals a smooth crossover from $d_{f}=2$ in a clean system to $d_{f}=1$ in a strongly disordered system, confirming the emergence of fractality and increased irregularity in the interface due to disorder.
We also calculated the density fluctuations in the system $\Delta\rho_{n}$. The detail of the calculations is provided in the Appendix. In Fig. 7(d) we show the plot of $\Delta\rho_{n}$ vs. $\rho_{n}$ for different $\rho_{d}=0.0$, $0.05$ , 0.1, 0.3 and $0.5$. For $\rho_{d}\leq 0.1$, the density fluctuations is large Dey et al. (2012) and $\Delta\rho_{n}$ goes as $\rho_{n}$, whereas for large $\rho_{d}$, $\Delta\rho_{n}$ $\sim$ $\sqrt{\rho_{n}}$ for large $\rho_{n}$. That suggests the diffusive nature of the particles for large disorder in the system.
IV Discussion
Disorder plays an important role due to its inherent intrinsic and extrinsic presence in natural systems, making its study highly relevant. Recently, research on disordered active matter has gained significant attention, emerging as an important area of study. In this work, we proposed a coarse-grained model of a collection of self-propelled particles to explore the phase behavior in the presence of pinned disorder. We use the disorder as pinning sites where the velocity of the particles become zero when it comes in contact to it. The key finding of our work is, weaker level of disorder promotes phase separation in the system within the phase space of $v_{0}$ and $\rho$ at lower values i.e. below the threshold values for $MIPS$ in a clean system. Additionally, the kinetics slows down at weak disorder, however beyond a certain threshold the system’s kinetic becomes arrested and phase separation suppresses.
Along with these key results, we conduct a detailed investigation of disordered system. Our analytical calculation using linear stability analysis shows a good agreement with the numerically obtained phase boundary at lower disorder level. We report that transition from the non-phase separated state to the phase-separated state in the system exhibits a crossover from discontinuous to continuous behavior with respect to $v_{0}$ as the pinning sites in the system increases. By analyzing the characteristic length scales and effective exponents, we quantify how the domain growth deviates from conventional LS law beyond a threshold of disorder density indicating the suppression of phase separation and kinetic arrest at higher disorder. The tail of structure factor fits well with the Porod’s tail for low disorder densities, but gradually deviates beyond a critical threshold, indicating increased interface roughness. This signature is observed in the cusp exponent which exhibits a crossover from 1 to 0.5 roughness of interface in the disordered system. Additionally, we calculate the fractal dimension of interface as a measure of it’s irregularity and find that $d_{f}$ varies from 2 to 1 with disorder. Further, we report the density fluctuations become giant in clean as well as weak disordered system and suppresses in systems with strong disorder. All the results combining in a bigger frame shows how disorder affects the system’s phase behavior, kinetics and dynamical properties.
The study of systems with disorder is very important as it allows researchers to explore how complex, real-world environments influence the dynamics and statistics of active systems. Disorder can arise in various forms, such as spatial heterogeneity, temporal fluctuations, or randomness in particle properties. Many natural systems, such as bacterial colonies, cell tissues, and animal groups, operate in environments that are inherently disordered. In this work, we provide a fundamental study on active matter system’s phase behavior with spatial heterogeneity as pinned disorder and can be helpful for understanding the behavior of biological systems in presence of disorder.
In this study, we explore the system with quenched obstacles, in future it would be interesting to explore with diffusive obstacles.
Acknowledgements.
P.J. gratefully acknowledge the DST INSPIRE fellowship for funding this project. The support and the resources provided by PARAM Shivay Facility under the National Supercomputing Mission, Government of India at the Indian Institute of Technology, Varanasi are gratefully acknowledged by all authors. S.M. thanks DST-SERB India, ECR/2017/000659, CRG/2021/006945 and MTR/2021/000438 for financial support. P.J. and S.M. also thank the Centre for Computing and Information Services at IIT (BHU), Varanasi.
V Data availability
The datasets used and/or analysed during the current study available from the corresponding author on reasonable request.
Appendix A Linearise calculation
A.1 Shifting of phase boundary
We develop a linearised calculation of the hydrodynamic equations provided in the main text and show the shifting of phase boundary in the plane of ($\rho_{0}$ and $v_{0}$). The homogeneous steady state solution of the above equations are $\bm{P}$=00 and $\rho$=$\rho_{0}$. Adding small fluctuations to the steady state solutions i.e; $\bm{P}=\delta\bm{P}$ and $\rho=\rho_{0}+\delta{\rho}$ the effective self-propulsion speed $v_{e}(\rho,\rho_{d})=v_{0}(1-\lambda\rho)A(r)$, where $\langle A(r)\rangle=1-\rho_{d}$, and the mean .$\langle$.$\rangle$ . . ⟩ is over different realizations. $\langle A(r)\rangle=1$ for the clean system. Further substituting the density in terms of small fluctuation $\delta\rho$, $v_{e}=v_{0}(1-\lambda\rho_{0})\langle A(r)\rangle-v_{0}\lambda\delta\rho\langle A(r)\rangle$. Now, we define $V=v_{0}(1-\lambda\rho_{0})\langle A(r)\rangle$. Equation 1 and 2 (main) becomes,
$\partial_{t}\delta\rho=-\bm{\nabla}.V\delta\bm{P}+D_{\rho}\bm{\nabla}^{2}\delta\rho-\bm{\nabla}.\bm{f_{\rho}}$ (3)
and
$\partial_{t}\delta\bm{P}=-\nu_{r}\delta\bm{P}-\frac{1}{2}V\delta\rho+\frac{1}{2}v_{0}\lambda\rho_{0}\langle A(r)\rangle\nabla\delta\rho+k\bm{\nabla}^{2}\delta\bm{P}+\bm{f_{P}}$ (4)
Taking divergence of equation(5) and defining .$\nabla.\delta\bm{P}=\theta$
$\partial_{t}\theta=-\nu_{r}\theta-\frac{1}{2}V\nabla^{2}\delta\rho+\frac{1}{2}v_{0}\lambda\rho_{0}\langle A(r)\rangle\nabla^{2}\delta\rho+k\bm{\nabla}^{2}\theta+\bm{\nabla}.\bm{f_{P}}$ (5)
Equation 3 becomes,
$\partial_{t}\delta\rho=-V\theta+D_{\rho}\bm{\nabla}^{2}\delta\rho-\bm{\nabla}.\bm{f}_{\rho}$ (6)
Later we replace the notation $\langle A(r)\rangle$ by $A$ for simplicity. To perform the mode analysis we write the Linearized equations for order-parameter and for density in Fourier mode,
$\theta(\bm{r},t)=\int_{\bm{q},\omega}\Theta(\bm{q},\omega)\exp(-i\omega t+i\bm{q}.\bm{r})d\bm{q}d\omega$
$\delta\rho(\bm{r},t)=\int_{\bm{q},\omega}\delta\rho(\bm{q},\omega)\exp(-i\omega t+i\bm{q}.\bm{r})d\bm{q}d\omega$
The linear equations in Fourier mode are,
$[-i\omega+D_{\rho}q^{2}]\delta\rho+{V}\Theta=-i\bm{q}.\bm{f}_{\rho}$ (7)
$[-\frac{1}{2}{V}+\frac{1}{2}v_{0}\rho_{0}\lambda A]q^{2}\delta\rho+[-i\omega+kq^{2}+\nu_{r}]\Theta=i\bm{q}.\bm{f_{P}}$ (8)
The linear equations can be easily solved for $\Theta$ and $\delta\rho$ using Matrix method. This can be written in 2 $\times$ 2 matrix form.
$\begin{bmatrix}-i\omega+D_{\rho}q^{2}&{V}\\ -\Omega q^{2}&-i\omega+kq^{2}+\nu_{r}\end{bmatrix}\begin{bmatrix}\delta\rho\\ \Theta\end{bmatrix}=\begin{bmatrix}-i\bm{q}.\bm{f_{\rho}}\\ i\bm{q}.\bm{f_{{P}}}\end{bmatrix}$
where $\Omega=\frac{1}{2}[V-v_{0}\rho_{0}\lambda A]$. The solutions for $\Theta$ and $\delta\rho$ is;
$\begin{bmatrix}\delta\rho\\ \Theta\par\end{bmatrix}=M^{-1}\begin{bmatrix}-iqf_{\rho}\\ iqf_{p}\end{bmatrix}$ (9)
Where the matrix M is,
$\begin{bmatrix}M\end{bmatrix}=\begin{bmatrix}-i\omega+D_{\rho}q^{2}&V\\ -\Omega q^{2}&-i\omega+kq^{2}+\nu_{r}\end{bmatrix}$
The inverse of 2 $\times$ 2 matrix is
$\begin{bmatrix}M^{-1}\end{bmatrix}=\frac{1}{det(M)}\begin{bmatrix}-i\omega+kq^{2}+\nu_{r}&-V\\ \Omega q^{2}&-\iota\omega+D_{\rho}q^{2}\end{bmatrix}$
where
$det[M]=-\omega^{2}-i\omega[(kq^{2}+\nu_{r})+D_{\rho}q^{2}]+D_{\rho}q^{2}(kq^{2}+\nu_{r})+V\Omega q^{2}$
The two modes obtained are
$w_{1}=-i\frac{1}{2}[(D_{\rho}+k)q^{2}+\nu_{r}]-i\frac{1}{2}[((\nu_{r}+kq^{2})-D_{\rho}q^{2})^{2}-4V\Omega q^{2}]^{1/2}$ (10)
$w_{2}=-i\frac{1}{2}[(D_{\rho}+k)q^{2}+\nu_{r}]+i\frac{1}{2}[((\nu_{r}+kq^{2})-D_{\rho}q^{2})^{2}-4V\Omega q^{2}]^{1/2}$ (11)
We can consider fluctuation as a wave has the form $e^{-i\omega t}$. This can be written as $e^{-iRe(w)}.e^{Im(w)}$. For instability $Im(w)>0$. Below we determine the condition for instability in the mode $\omega_{2}$,
$((\nu_{r}+kq^{2})-D_{\rho}q^{2})^{2}-4V\Omega q^{2}>0$
$\sqrt{((\nu_{r}+kq^{2})-D_{\rho}q^{2})^{2}-4V\Omega q^{2}}>(D_{\rho}+k)q^{2}+\nu_{r}$
$((\nu_{r}+kq^{2})-D_{\rho}q^{2})^{2}-4V\Omega q^{2}>[(D_{\rho}+k)q^{2}+\nu_{r}]^{2}$
$(\nu_{r}+kq^{2})^{2}+(D_{\rho}q^{2})^{2}-2D_{\rho}q^{2}(\nu_{r}+kq^{2})-4V\Omega q^{2}-(D_{\rho}+k)^{2}q^{4}-\nu_{r}^{2}-2\nu_{r}q^{2}(D_{\rho}+k)>0$
$-4q^{2}[D_{\rho}(kq^{2}+\nu_{r})+V\Omega]>0$
To satisfy this, $V\Omega<0$ and $|V\Omega|>D_{\rho}(kq^{2}+\nu_{r})$.
case-1 $V\Omega<0$
$\frac{1}{2}[v_{0}^{2}A^{2}(1-\lambda\rho_{0})(1-2\lambda\rho_{0})]<0$
$(1-\lambda\rho_{0})(1-2\lambda\rho_{0})<0$
case-2 $|V\Omega|>D_{\rho}(kq^{2}+\nu_{r})$
$\frac{1}{2}[v_{0}^{2}A^{2}(1-\lambda\rho_{0})(1-2\lambda\rho_{0})]>D_{\rho}(kq^{2}+\nu_{r})$
$(2\lambda^{2}\rho_{0}^{2}-3\lambda\rho_{0}+1)>\frac{2D_{\rho}(kq^{2}+\nu_{r})}{v_{0}^{2}A^{2}}$
for small q,
$02\lambda^{2}\rho_{0}^{2}+1-3\lambda\rho_{0}-\frac{2D_{\rho}\nu_{r}}{v_{0}^{2}A^{2}}<$ 0
$\rho_{0}^{2}-\frac{3\rho_{0}}{2\lambda}+\frac{1}{2\lambda^{2}}-\frac{D_{\rho}\nu_{r}}{\lambda^{2}v_{0}^{2}A^{2}}<0$
solving this quadratic equation for $\rho_{0}$,
$\rho_{0\pm}=\frac{3}{4\lambda}\pm\frac{1}{4\lambda}\sqrt{1+\frac{16D_{\rho}\nu_{r}}{v_{0}^{2}A^{2}}}$
Rewriting the equations when the dimensionless activity $\frac{v_{0}^{2}}{D_{\rho}\nu_{r}}$ is larger and rearranging the terms we can write,
$\rho_{0+}v_{0}^{2}=A_{1}v_{0}^{2}+\frac{B_{1}}{A^{2}}$ (12)
$\rho_{0-}v_{0}^{2}=A_{2}v_{0}^{2}-\frac{B_{2}}{A^{2}}$ (13)
here, $A_{1},A_{2},B_{1},B_{2}$ are parameter dependent constants. For the clean system $A=1$. Out of the two solutions as given in Eqs. 12 and 13, we first analyze the second solution Eq. 13 and in Fig. 3(a) in the manuscript we have shown the straight line plot between $\rho_{0}v_{0}^{2}$ vs. $v_{0}^{2}$. The region above the straight line drawn with solid line shows the instability or the phase separation. However, in the presence of disorder $A=(1-\rho_{d})$, the straight line between $\rho v_{0}^{2}$ vs. $v_{0}^{2}$ shifts down and additional region marked with shaded color shows the instability. Which makes the mean density $\rho_{0}$ and activity $v_{0}$ shifts towards the smaller values for phase separation in the presence of disorder.
A.2 Calculation of Structure factor
From the Eq. 9, the solution for $\delta\rho$ is;
$\delta\rho(\bm{q},\omega)=\frac{[-(-i\omega+kq^{2}+\nu_{r})iqf_{q}-iqf_{P}V]}{(\omega-i\omega_{1})(\omega-i\omega_{2})}$ (14)
The correlation function is defined as $C_{\rho\rho}(\bm{q},\omega)$ = $\langle\delta\rho(\bm{q},\omega)\delta\rho(\bm{-q},-\omega)\rangle$;
$\langle\delta\rho(\bm{q},\omega)\delta\rho(\bm{-q},-\omega)\rangle=\frac{[\omega^{2}+(kq^{2}+\nu_{r})^{2}]q^{2}\Delta_{\rho}+V^{2}q^{2}\Delta_{P}}{(\omega^{2}+\omega_{1}^{2})(\omega^{2}+\omega_{2}^{2})}$ (15)
The static structure factor is defined as,
[…]
. The full expression for the static structure factor is;
$S(q)=\frac{q^{2}\Delta_{\rho}\pi}{(D_{q}+k)q^{2}+\nu_{r}}+\frac{\pi v_{0}^{2}(1-\lambda\rho_{0})^{2}A^{2}\Delta_{P}}{4[(kq^{2}+\nu_{r})D_{\rho}+\frac{A^{2}}{2}v_{0}^{2}(1-\lambda\rho_{0})(1-2\rho_{0}\lambda)][(D_{\rho}+k)q^{2}+\nu_{r}]}$ (16)
For clean system limit, setting $A=1$ we can get,
$S(q)=\frac{q^{2}\Delta_{\rho}\pi}{(D_{q}+k)q^{2}+\nu_{r}}+\frac{\pi v_{0}^{2}(1-\lambda\rho_{0})^{2}\Delta_{P}}{4[(kq^{2}+\nu_{r})D_{\rho}+\frac{v_{0}^{2}}{2}(1-\lambda\rho_{0})(1-2\rho_{0}\lambda)]+(D_{\rho}+k)q^{2}+\nu_{r})}$ (17)
which matches well with the result obtained in Fily and Marchetti (2012). In the presence of disorder ($A=1-\rho_{d}$), the second term in eq. 16 flatten the structure factor. This particular behavior is also observed in Fig. 7(a), where for clean system and weak disorder, the decay is sharp but for higher $\rho_{d}$ the plot becomes flatter and the tail of structure factor deviates from Porod’s tail.
A.3 Fractal dimension of interface
To investigate the fractal nature of the domains we plot the fractal dimension of the interface $d_{f}$ vs. disorder $\rho_{d}$ in Fig.8 (c) (main manuscript). To calculate the fractal dimension of the interface we follows the below procedure: For a fractal interface, the length of the interface decreases with time $t$ as $B(t)\approx t^{-\frac{d_{f}-1}{3}}$, where, $d_{f}$ is the fractal dimension Streitenberger (2000). The length of the interface is determined using a box-counting algorithm. We begin by evaluating the fluctuation of the order parameter at each lattice point $(i,j)$ as $\delta\rho(i,j)=\rho(i,j)-\rho_{0}$ where, $\rho_{0}$ is the mean value of $\rho$. A point is classified as part of a high-density region if $\delta\rho>0$, otherwise, it is considered to belong to a low-density region. Each lattice point is then further categorized based on the number of its high-density neighbors, denoted by by $n_{h}$. Points with $n_{h}=4$ are considered core points, residing within clusters. Points with $2\leq n_{h}\leq 3$ that has at least one neighbor which is a core point are classified as edge points, lying at the interface. Points outside these criteria are deemed isolated, located in low-density regions. In this study, the spatial grid size is uniform in both $x$ and $y$-directions. The perimeter or length of the interface, $B(t)$, is then calculated by multiplying the number of edge points by the spatial grid size.
A.4 Density fluctuation
To calculate the density fluctuations in the system $\Delta\rho_{n}$, we divide the simulation box of size $512\times 512$ into square blocks of different sizes from $1\times 1$ to $256\times 256$. For each block size, we calculate the total density within each block across multiple time steps in the steady state and determine the average density $\rho_{n}$, so as the fluctuates from that average. By repeating this process for different block sizes, we calculate average densities $\rho_{n}$ and their variances $\Delta\rho_{n}$ for different system with $\rho_{d}=0-0.5$